OpenStudy (anonymous):

How do you calculate force normal and force friction if the object has a constant acceleration?

OpenStudy (anonymous):

It depends on what type of platform the object is on.

OpenStudy (anonymous):

 Take the instance of an object on a flat surface. In this case, we can express the sum of the forces in the y-direction as the following \[\sum F_y=N-mg=0 \rightarrow N=mg\] This means that the normal force exerted on the object by the surface is simply that objects weight! Also, friction is always modeled as:\[F_f=\mu N\] Where μ is the coefficient of friction. In this scenario, since N is just mg, then our equation for frictional force is: \[F_f=\mu mg\]

OpenStudy (anonymous):